OpenStudy (anonymous):

A jogger started a course and jogged an average speed of 4mph. One hour later, a cyclist started the same course and cycled at an average speed of 11mph. How long after the jogger started did the cyclist overtake the jogger?

OpenStudy (mattfeury):

So you can set up an algebraic equation to solve this. You basically want to know the time at which the jogger and the cyclist are at the same position. so we have an unknown variable time, which we call 't'

OpenStudy (mattfeury):

for the jogger, 4mph * t = distance travelled for the cyclist, 11mph * (t-1) = distance travelled. this is (t-1) because when t = 1, the cyclist is just beginning.

OpenStudy (mattfeury):

\[4*t = 11*(t-1)\]\[4*t = 11*t - 11\]\[11 = 7*t\]\[t = \frac{11}{7} seconds\]